Let f(x) = x2 - 20x + 100. Find x so that f(x) = 64
I am completely lost. This is what I did so far. But I know i am wrong.
a=1 b=-20 x=100
x= {-20 +/- sq rt (400-400)}/2
x= {-20 +/- sq rt 0}/2
x= 10

Answer by checkley77(12844)

Let f(x) = x2 - 20x + 100
f(64)=64^2-20*64+100
F(64)=4,096-1,280+100
F(64)=2,916 ANS.